# CastMind: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting

### v2

@@ 4.4 Exploration Analysis
Figure 6: Feature-level and model-level case analysis of CastMind across datasets. The heatmaps report the relative usage frequency (%) of temporal features and candidate forecasting models, highlighting a preference for statistical features and deep learning models.
Figure 7: Case-study comparisons on four datasets (ETTm, MOPEX, Windy Power, and DE), where CastMind produces predictions that track the ground truth more closely than the cased-based prediction, especially under periodic or highly volatile patterns.
@@
To investigate the impact of the foundation LLM on predictive performance, we evaluate the framework using diverse LLM backbones, including DeepSeek-V3 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), GPT-4o (Achiam et al., 2023), Gemini-2.5-pro (Comanici et al., 2025), and GPT-5 (Singh et al., 2025). As illustrated in Figure 4, GPT-5 consistently achieves the lowest MAE across diverse datasets such as Windy Power, DE, and PJM, establishing a clear performance advantage over other LLMs. The superior performance of GPT-5 is attributed to its powerful pre-training and reasoning capabilities, enabling it to effectively handle complex time series, adapt to various context, and provide more accurate predictions. Notably, the performance gap becomes most distinct on the Windy Power dataset, where GPT-5 demonstrates a significantly stronger ability to model complex environmental dependencies and high-frequency volatility compared to competitive baselines.
@@ 4.5 Case Studies
We analyze the usage frequency within the feature set and case library to understand CastMind’s preferences. As shown in Figure 6, heatmaps reveal selection patterns driven by data characteristics. In the feature domain, CastMind favors fundamental statistical features, suggesting that reliable global properties serve as better foundations than volatile complexity measures. Meanwhile, model selection shows a clear bias towards deep learning architectures, with Transformers consistently dominating the candidate pool over statistical models, confirming that high-capacity backbones provide the necessary prior. Ultimately, the varying distribution across datasets confirms that the framework dynamically adjusts its retrieval strategy, rather than relying on static priors.
@@
To assess forecasting quality, we compare final predictions with the case-based prediction and ground truth. As shown in Figure 7, the case-based prediction often suffers from over-smoothing, generating conservative trajectories that miss rapid fluctuations. This is evident on MOPEX and Windy Power, where the candidate models misses significant peaks and reverts to the mean. In contrast, CastMind shows superior sensitivity, closely tracking the ground truth even under high volatility on the ETTm and DE. This confirms that reasoning allows the model to escape the mean-reversion trap, ensuring forecasts retain the original series’ characteristics. A detailed case study is provided in the AppendixB.4.
@@ 5 Conclusion
In this work, we introduces CastMind, an interaction-driven agentic reasoning framework designed to bridge the gap between existing time series forecasting models and expert-driven iterative reasoning. CastMind redefines the forecasting process as a multi-turn iteration, integrating context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn autonomous interaction process. The proposed framework leverages training-free LLMs as reasoning engines, complemented by a lightweight toolkit that effectively combines human wisdom with LLM intelligence. Comprehensive evaluations demonstrate that CastMind’s ability to model expert-like iterative reasoning and refinement is key to its superior performance in both short-term and long-term forecasting tasks. Although the results are preliminary, our findings suggest that forecasting is shifting from a static regression-based approach to a interaction-driven agentic reasoning paradigm.
@@
This work advances time series forecasting through CastMind, an interaction-driven agentic framework that enhances transparency and decision reliability by mimicking expert-like iterative reasoning. By transforming forecasting into a multi-stage, autonomous interaction process supported by external context, our approach significantly improves interpretability in critical sectors such as infrastructure management and financial planning. However, we acknowledge that the multi-turn interaction with LLMs incurs higher inference latency and energy consumption compared to static, single-pass regression models. Consequently, we emphasize that deployment in high-stakes environments should remain subject to rigorous human oversight to mitigate potential risks associated with autonomous agentic decision-making. The information in the contextual pool and knowledge base is not directly sourced from the dataset itself, thus there is no risk of data leakage.
@@ A.1 Feature Overview
Table 5: Mathematical definitions of the twenty extracted temporal features used in CastMind, covering distributional statistics, temporal dependence, spectral complexity, structural heterogeneity, and seasonality, serving as a compact summary of key temporal characteristics for forecasting.
@@ B.3 Complete Experiment
Table 7: Forecasting performance on short-term and long-term benchmarks. We compare CastMind with each individual candidate model in the case library to show the benefit of interaction-driven reasoning and iterative refinement.

### v3

@@ 4.4 Exploration Analysis
Figure 6: Feature-level and model-level case analysis of AlphaCast across datasets. The heatmaps report the relative usage frequency (%) of temporal features and candidate forecasting models, highlighting a preference for statistical features and deep learning models.
Figure 7: Case-study comparisons on four datasets (ETTm, MOPEX, Windy Power, and DE), where AlphaCast produces predictions that track the ground truth more closely than the cased-based prediction, especially under periodic or highly volatile patterns.
@@
To investigate the impact of the foundation LLM on predictive performance, we evaluate the framework using diverse LLM backbones, including DeepSeek-V3 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), GPT-4o (Achiam et al., 2023), Gemini-2.5-pro (Comanici et al., 2025), and GPT-5 (Singh et al., 2025). As illustrated in Figure 4, GPT-5 consistently achieves the lowest MAE across diverse datasets such as Windy Power, DE, and PJM, establishing a clear performance advantage over other LLMs. The superior performance of GPT-5 is attributed to its powerful pre-training and reasoning capabilities, enabling it to effectively handle complex time series, adapt to various contexts, and provide more accurate predictions. Notably, the performance gap becomes most distinct on the Windy Power dataset, where GPT-5 demonstrates a significantly stronger ability to model complex environmental dependencies and high-frequency volatility compared to competitive baselines.
@@ 4.5 Case Studies
We analyze the usage frequency within the feature set and case library to understand AlphaCast’s preferences. As shown in Figure 6, heatmaps reveal selection patterns driven by data characteristics. In the feature domain, AlphaCast favors fundamental statistical features, suggesting that reliable global properties serve as better foundations than volatile complexity measures. Meanwhile, model selection shows a clear bias towards deep learning architectures, with Transformers consistently dominating the candidate pool over statistical models, confirming that high-capacity backbones provide the necessary prior. Ultimately, the varying distribution across datasets confirms that the framework dynamically adjusts its retrieval strategy, rather than relying on priors.
@@
To assess forecasting quality, we compare final predictions with the case-based prediction and ground truth. As shown in Figure 7, the case-based prediction often suffers from over-smoothing, generating conservative trajectories that miss rapid fluctuations. This is evident on MOPEX and Windy Power, where the candidate models miss significant peaks and reverts to the mean. In contrast, AlphaCast shows superior sensitivity, closely tracking the ground truth even under high volatility on the ETTm and DE. This confirms that reasoning allows the model to escape the mean-reversion trap, ensuring forecasts retain the original series’ characteristics. A detailed case study is provided in the AppendixB.4.
@@ 5 Conclusion
In this work, we introduce AlphaCast, an interaction-driven agentic reasoning framework designed to bridge the gap between existing time series forecasting models and expert-driven iterative reasoning. AlphaCast redefines the forecasting process as a multi-turn iteration, integrating context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn autonomous interaction process. The proposed framework leverages training-free LLMs as reasoning engines, complemented by a lightweight toolkit that effectively combines human wisdom with LLM intelligence. Comprehensive evaluations demonstrate that AlphaCast’s ability to model expert-like iterative reasoning and refinement is key to its superior performance in both short-term and long-term forecasting tasks. Although the results are preliminary, our findings suggest that forecasting is shifting from a static regression-based approach to a interaction-driven agentic reasoning paradigm.
@@
This work advances time series forecasting through AlphaCast, an interaction-driven agentic framework that enhances transparency and decision reliability by mimicking expert-like iterative reasoning. By transforming forecasting into a multi-stage, autonomous interaction process supported by external context, our approach significantly improves interpretability in critical sectors such as infrastructure management and financial planning. However, we acknowledge that the multi-turn interaction with LLMs incurs higher inference latency and energy consumption compared to static, single-pass regression models. Consequently, we emphasize that deployment in high-stakes environments should remain subject to rigorous human oversight to mitigate potential risks associated with autonomous agentic decision-making. The information in the contextual pool and knowledge base is not directly sourced from the dataset itself, thus there is no risk of data leakage.
@@ A.1 Feature Overview
Table 5: Mathematical definitions of the twenty extracted temporal features used in AlphaCast, covering distributional statistics, temporal dependence, spectral complexity, structural heterogeneity, and seasonality, serving as a compact summary of key temporal characteristics for forecasting.
@@ B.3 Complete Experiment
Table 7: Forecasting performance on short-term and long-term benchmarks. We compare AlphaCast with each individual candidate model in the case library to show the benefit of interaction-driven reasoning and iterative refinement.